OpenStudy (anonymous):

m= -1;(-3,6) write a point-slope equation for the line with the given slope that contains the given point

OpenStudy (anonymous):

are you on algebra 1 or 2?

OpenStudy (anonymous):

do you have a y- intercept?

OpenStudy (anonymous):

no i dont have the Y intercept

OpenStudy (anonymous):

so m=-1 and b=6

OpenStudy (goformit100):



OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

now what you have to do is make an equation: y=mx+b so y=-1x+6 now make x= 0 and find y and then substitute y in and you get your answer! do you understand?

OpenStudy (anonymous):

ok so i would write y=mx+b y=-1x+6 y=-1(0)+6 y=6 is this write

OpenStudy (anonymous):

yes! now pick 2 other x numbers to get more y

OpenStudy (anonymous):

now thats what i get confused on

OpenStudy (anonymous):

well actually on this problem you dont need to do that ... my mistake... all you have to do is write out your equation which is...

OpenStudy (anonymous):

if you have a point and a slope use the formula y - y1 = m ( x - x1) where x1 and y1 are the points given to you and m slope so y - 6 = -1 [ x - (-3) ] now just solvev for y

OpenStudy (anonymous):

how do I know what two number im looking for? is that when i plug in the numbers i have already with a formula? im not sure if im even asking this question right sorry

OpenStudy (anonymous):

you can actually pick any numbers that you want

OpenStudy (anonymous):

plug in the coordinates you have, so x coordinate is x1 and your y coordinate is y1.

OpenStudy (anonymous):

your ordered pair that you have in the equation.

OpenStudy (anonymous):

poetrylover i dont think thats right

OpenStudy (anonymous):

since your original equation isnt in correct format you would have to change it to y=mx+b.

OpenStudy (anonymous):

i thought thats what i did

OpenStudy (anonymous):

yeah it is but do you know wht the final equation is?

OpenStudy (anonymous):

no I feel confused

OpenStudy (anonymous):

so y = -x - 3 + 6 => y = -x + 3 lets check the point x = -3, the y has to be 6 in order to be true so y = -(-3) + 3 = 3 + 3 = 6

OpenStudy (anonymous):

okay hold on. y - 6 = -1(x - (-3))

OpenStudy (anonymous):

this is how you write it in point slope form.

OpenStudy (anonymous):

y - 6 = -1(x+3) y-6 =-x -3 +6 +6 y = -x + 3 this is your answer

OpenStudy (anonymous):